# GPU-Accelerated Gravitational Lensing &\&& Dynamical (GLaD) Modeling for Cosmology and Galaxies

### v1

@@ 6 Summary and outlook
In this paper, we present a GPU-accelerated code (GLaD) for self-consistent lensing and dynamical modeling, based on Yıldırım et al. (2020) for the lensing part and on Cappellari (2020) for the dynamics part. This method combines lensing and dynamical models by solving the Jeans equations in an axisymmetric geometry. The primary purpose of this code is for time-delay cosmography, but it can also be naturally applied to galaxy evolution studies (Shajib et al. 2021; Tan et al. 2024; Sheu et al. 2024; Sahu et al. 2024).
@@
We simulate the lensing and kinematic data for the lensed quasar system RXJ1131 to test whether GLaD can recover the mock input value. Since the lens galaxy in RXJ1131 exhibits a central velocity dispersion $\geq 300\rm\leavevmode\nobreak\ km\leavevmode\nobreak\ s^{-1}$, we add a $M_{\rm BH}=5\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ in the mock mass profile. For the kinematic map, we generate one ideal kinematic map with the $2\%$ statistical error and a biased kinematic map with a $5\%$ systematic bias (as a worst-case scenario) in all the velocities. We use GLaD to perform the joint modeling on the simulated data to test the influence of the BH and the systematic error in the kinematic map. We found as follows:
@@
  • Given the high BH mass of $5\times 10^{9}\ {\rm M_{\odot}}$ adopted in our mock data, the BH mass plays a crucial role in constraining $\beta_{\rm ani}$ and $D_{\rm d}$. By adjusting $\beta_{\rm ani}$, one can mimic the effect of a massive BH, making it difficult to constrain anisotropy without precise knowledge of the BH mass. Additionally, $\beta_{\rm ani}$ is positively correlated with $D_{\rm d}$, meaning any bias in the inferred $\beta_{\rm ani}$ leads to a corresponding bias in $D_{\rm d}$. As shown in Sect. 5.2, modeling with an incorrect BH mass results in an inferred $H_{0}$ that is 3% higher than the mock input value.
  • In Sect. 5.2, we present two approaches to mitigate the impact of the BH on $H_{0}$ measurements. The first approach involves using insights from nearby galaxies to determine the most probable range for the BH mass. We then perform a series of models with BH mass variations within this range, combining the results using the BIC weights to obtain an unbiased distance and $H_{0}$ measurement. The advantage of this approach is that it leverages the full kinematic dataset and a well-motivated prior. However, the disadvantage is the need to run multiple models. In the second approach, we bypass the sensitivity of the kinematic data to the BH mass by excluding the central kinematic bins, allowing us to retrieve the $H_{0}$ value with just one model, without significant reliance on prior knowledge.
  • The systematic bias in spatially resolved kinematic data does not impact the constraints on $\lambda_{\rm int}$ and $D_{\rm\Delta t,int}$, as these parameters are influenced by the shape of the 2D $v_{\rm rms}$ distribution. However, an overall bias in the kinematic data does not alter the shape of $v_{\rm rms}$; it only affects its amplitude.
GLaD will be applied to the NIRSpec IFU observations of the lens galaxy in RXJ1131. Using simulated data, we identified a trade-off between the BH mass and the anisotropy parameter $\beta_{\rm ani}$, as well as the influence of BH mass on $D_{\rm d}$ in this paper. In our simulated kinematic dataset, we used a higher BH mass compared to the value from the $M_{\rm BH}-\sigma_{\rm disp}$ relation. We aim to determine whether these effects are also present in real observations. If confirmed, we can further explore strategies to mitigate potential biases in $D_{\rm d}$.
As for the second test in this paper on systematic bias in the kinematic data, its impact on future $H_{0}$ measurements is expected to be minor, given the recent work by Knabel et al. (2025) who demonstrated that systematics errors of kinematic measurements can be controlled at the sub-percent level.
@@
Our study highlights the speed gains achieved by using a single GPU, and in the future, parallelizing computations across multiple GPUs could further improve efficiency. Our developments will enable more efficient lensing and dynamical modeling of galaxies with high quality data for future cosmological and galaxy studies.
@@ Acknowledgements
We thank Tommaso Treu, Shawn Knabel, Simon Birrer and Xiang-Yu Huang for helpful discussions and feedback on this work. HW and SHS thank the Max Planck Society for support through the Max Planck Fellowship for SHS. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LENSNOVA: grant agreement No 771776). This work is supported in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311. AG acknowledges the Swiss National Science Foundation (SNSF) for supporting this work.
Appendix A Implementation of the enfw profile
@@
where $r_{\rm s}$ is the scale radius and $\rho_{\rm s}$ is the characteristic density. In general, Eq. 62 does not yield an analytical expressions for lensing properties. Instead, computationally demanding numerical integration has to be performed. The idea in Oguri (2021) is to decompose the Eq. 62 into a series of basis functions, i.e., core steep ellipsoids (CSEs) which has simple analytical expressions of SL properties such as deflection angles $\bm{\alpha}_{\rm enfw}$ and the lensing potential $\psi_{\rm enfw}$.
@@
they achieved an accuracy of $\sigma=10^{-4}$ in recovering $\kappa_{\rm gNFW}$ using CSEs, with $u_{j}$ spanning a wide range from $10^{-6}$ to $10^{3}$. The amplitude $A_{i}$ and core radius $s_{i}$ are predetermined before evaluating the lensing properties of $\kappa_{\rm gNFW}$ for any given values of $\rho_{s}$, $r_{s}$, and $q$ [^14] Note that $\rho_{s}$ is omitted in Eq. 66 because it acts as a constant scaling factor and does not affect the decomposition process. The corresponding lens potential of individual CSE is
@@
Appendix B Implementation of the EPL profile
@@
Appendix C Joint modeling with ideal kinematic data across varying source grid resolutions
@@
Appendix D Joint modeling using kinematics data exclude the central bins
Figure 15: The comparison between the joint modeling using the full ideal kinematic data (orange contours) and the ideal kinematic data (green contours) excluding the central regions. We adopt no BH in the composite mass model. We observe that uncertainties of parameters are enlarged but the measurements of $D_{\rm d}$ and $\beta_{\rm ani}$ move toward the mock input values using the kinematic map excluding the central regions.
Figure 16: The comparison between the joint modeling using the full ideal kinematic data (orange contours) and the ideal kinematic data (green contours) excluding the central regions. We adopt $M_{\rm BH}=7\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ in the composite mass model.
Appendix E The BIC weight factor $f_{\textnormal{BIC}}^{*}$ to joint models
Data Model Source Resolution $\chi^{2}_{\text{dyn}}$ $f^{*}_{\text{BIC}}$
@@ Appendix E The BIC weight factor $f_{\textnormal{BIC}}^{*}$ to joint models
COMPOSITE 68 54.11 0.1272
66 55.08 0.0785
64 54.61 0.0991
$M_{\text{BH}}=1\times 10^{9}M_{\odot}$ 62 54.21 0.1213
60 54.30 0.1160
Lensing & Dynamics IDEAL $M_{\text{BH}}=2\times 10^{9}M_{\odot}$ 68 50.61 0.7204
66 50.36 0.8171
64 50.37 0.8129
62 50.48 0.7712
60 50.45 0.7805
$M_{\text{BH}}=3\times 10^{9}M_{\odot}$ 68 50.26 0.8612
66 49.96 0.9799
64 50.00 0.9672
62 50.06 0.9482
60 50.03 0.9567
$M_{\text{BH}}=4\times 10^{9}M_{\odot}$ 68 50.34 0.8327
66 50.10 0.9325
64 50.15 0.9065
62 50.20 0.8866
60 50.16 0.9050
$M_{\text{BH}}=5\times 10^{9}M_{\odot}$ 68 50.87 0.6350
66 50.79 0.6581
64 50.73 0.6790
62 50.65 0.7080
60 50.75 0.6723
$M_{\text{BH}}=6\times 10^{9}M_{\odot}$ 68 51.71 0.4162
66 51.86 0.3865
64 51.81 0.3965
62 51.74 0.4094
60 51.67 0.4238
Data Model Source Resolution $\chi^{2}_{\text{dyn}}$ $f^{*}_{\text{BIC}}$
@@
COMPOSITE 68 52.98 0.2219
66 53.55 0.1667
64 53.28 0.1904
$M_{\text{BH}}=7\times 10^{9}M_{\odot}$ 62 53.24 0.1949
60 53.36 0.1835
Lensing & Dynamics IDEAL $M_{\text{BH}}=8\times 10^{9}M_{\odot}$ 68 54.85 0.0870
66 55.49 0.0633
64 55.11 0.0763
62 55.05 0.0787
60 55.13 0.0758
$M_{\text{BH}}=9\times 10^{9}M_{\odot}$ 68 56.95 0.0307
66 57.94 0.0187
64 57.21 0.0269
62 57.22 0.0268
60 57.53 0.0230
$M_{\text{BH}}=10\times 10^{9}M_{\odot}$ 68 60.93 0.0042
66 60.67 0.0047
64 60.73 0.0040
62 60.12 0.0052
60 60.45 0.0052
EPL Model 68 58.15 0.1338
66 58.12 0.1356
64 58.33 0.1227
62 60.56 0.0400
60 58.29 0.1247
Table 6: Continuation of the previous table: Comparison of models at different source resolutions, showing $\chi_{\text{dyn}}^{2}$ and $f_{\text{BIC}}^{*}$.

### v2

@@ 6 Summary and outlook
In this paper, we present a GPU-accelerated code (GLaD) for self-consistent lensing and dynamical modeling, based on Yıldırım et al. (2020) for the lensing part and on Cappellari (2020) for the dynamics part. This method combines lensing and dynamical models by solving the Jeans equations in an axisymmetric geometry. The primary purpose of this code is for time-delay cosmography, but it can also be naturally applied to galaxy evolution studies (Shajib et al. 2021; Tan et al. 2024; Sheu et al. 2025; Sahu et al. 2024).
@@
We simulate the lensing and kinematic data for the lensed quasar system RXJ1131 to test whether GLaD can recover the mock input value. Since the lens galaxy in RXJ1131 exhibits a central velocity dispersion $\geq 300\rm\leavevmode\nobreak\ km\leavevmode\nobreak\ s^{-1}$, we add a $M_{\rm BH}=5\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ in the mock mass profile. For the kinematic map, we generate one ideal kinematic map with the $2\%$ statistical error and a biased kinematic map with a $5\%$ systematic bias (as a worst-case scenario) in all the velocities. We use GLaD to perform the joint modeling on the simulated data to test the influence of the systematic error and BH effects in the kinematic map. We found as follows:
@@
  • Systematic biases in the overall amplitude of spatially resolved kinematic data do not significantly impact the constraints on $\lambda_{\rm int}$ and $D_{\rm\Delta t,int}$, because these parameters are primarily sensitive to the shape of the two-dimensional $v_{\rm rms}$ distribution. Since an overall bias alters only the amplitude and not the shape of $v_{\rm rms}$, its effect on the inference of these parameters is minimal.
@@
  • Given the high BH mass of $5\times 10^{9}\ {\rm M_{\odot}}$ adopted in our mock data, the BH mass plays a crucial role in constraining $\beta_{\rm ani}$ and $D_{\rm d}$. By adjusting $\beta_{\rm ani}$, one can mimic the effect of a massive BH, making it difficult to constrain anisotropy without precise knowledge of the BH mass. Additionally, $\beta_{\rm ani}$ is positively correlated with $D_{\rm d}$, meaning any bias in the inferred $\beta_{\rm ani}$ leads to a corresponding bias in $D_{\rm d}$. As shown in Sect. 5.4, modeling with an incorrect BH mass results in an inferred $H_{0}$ that is 3% higher than the mock input value.
  • In Sects. 5.4 and 5.5, we present two approaches to mitigate the impact of the BH on $H_{0}$ measurements. The first approach involves using insights from nearby galaxies to determine the most probable range for the BH mass. We then perform a series of models with BH mass variations within this range, combining the results using the BIC weights to obtain an unbiased distance and $H_{0}$ measurement. The advantage of this approach is that it leverages the full kinematic dataset and a well-motivated prior. However, the disadvantage is the need to run multiple models. In the second approach, we bypass the sensitivity of the kinematic data to the BH mass by excluding the central kinematic bins, allowing us to retrieve the $H_{0}$ value with just one model.
GLaD will be applied to the NIRSpec IFU observations of the lens galaxy in RXJ1131. Using the ideal kinematic simulated data, we identified a trade-off between the BH mass and the anisotropy parameter $\beta_{\rm ani}$, as well as the influence of BH mass on $D_{\rm d}$ in this paper. In our simulated kinematic dataset, we used a higher BH mass compared to the value from the $M_{\rm BH}-\sigma_{\rm disp}$ relation. We aim to determine whether these effects are also present in real observations. If confirmed, this would motivate further exploration of strategies to mitigate potential biases in the inference of $D_{\rm d}$. One promising approach would be to adopt a radially varying anisotropy profile, $\beta_{\rm ani}(r)$, with carefully chosen priors, as proposed in Simon et al. (2024). This would be especially helpful in tightening the constraints on $\beta_{\rm ani}$ values near the BH.
As for the test on systematic bias in the kinematic data, its impact on future $H_{0}$ measurements is expected to be minor, given the recent work by Knabel et al. (2025) who demonstrated that systematics errors of kinematic measurements can be controlled at the sub-percent level.
@@
Our study highlights the speed gains achieved by using a single GPU, and in the future, parallelizing computations across multiple GPUs could further improve efficiency. Our developments will enable more efficient lensing and dynamical modeling of galaxies with high-quality data for future cosmological and galaxy studies.
@@ Acknowledgements
We would like to thank the anonymous referee whose comments were helpful in improving the paper. We thank Tommaso Treu, Shawn Knabel, Simon Birrer and Xiang-Yu Huang for helpful discussions and feedback on this work.
HW and SHS thank the Max Planck Society for support through the Max Planck Fellowship for SHS. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LENSNOVA: grant agreement No 771776). This work is supported in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311.
AG acknowledges the Swiss National Science Foundation (SNSF) for supporting this work.
AJS received support from NASA through the STScI grants HST-GO-16773 and JWST-GO-2974.
Appendix A MGE to surface brightness and mass convergence
The observed 2D SB of the lens galaxies, $I(x^{\prime},y^{\prime})$  [^15] Note that we present the general case here. In this paper, we perform 1D MGE fitting (see Sect. 3.2) to model the profile along the radial direction $R$ with a fixed axis ratio $q$, i.e., $R=\sqrt{x^{2}+y^{2}/q^{2}}$., is expressed through multiple Gaussians:
The deprojection process depends on the assumption of the galaxies’ shapes. For the commonly found elliptical galaxies with an oblate shape, the deprojection requires:
$\cos^{2}{i}<q_{\rm min}^{\prime 2}$ (59)
where $r$ is the 3D radial distance to the galaxy centroid, $\theta$ is the polar angle (see definition in Cappellari 2020, Fig. 1), $\sigma_{j}=\sigma^{\prime}_{j}$ and $q_{j}=\frac{\sqrt{q^{\prime 2}_{j}-\cos{i}^{2}}}{\sin{i}}$ denote the dispersion and axis ratio of Gaussians after deprojection. The potential $\psi_{\rm D,int}$ in Eqs. 18 and 19 is derived by integrating the MGE of the 3D mass density $\rho_{\rm int}$. Following the approach used to infer the light tracer $\rho_{*}$, the 3D density profile $\rho_{\rm int}$ is obtained by deprojecting $\Sigma_{\rm int}$ (see Eq. 16). The surface mass density $\Sigma_{\rm int}$ is expressed as a sum of multiple Gaussian components,
Note we use index $i$ to denote the MGE components of the mass density, and $j$ for the luminosity components. The set of Gaussians describing the SB of lens galaxies (see Eqs. 58,  61) is not necessarily identical to the MGEs of their mass densities. Therefore, $i\neq j$ meaning that $\sigma^{\prime}_{i}\neq\sigma^{\prime}_{j}$, $q^{\prime}_{i}\neq q^{\prime}_{j}$ and $M\neq N$ unless mass follows light. The deprojected $\rho_{\rm int}(r,\theta)$ is
Appendix B Implementation of the enfw profile
@@
where $r_{\rm s}$ is the scale radius and $\rho_{\rm s}$ is the characteristic density. In general, Eq. 63 does not yield analytical expressions for lensing properties. Instead, computationally demanding numerical integration has to be performed. The idea in Oguri (2021) is to decompose the Eq. 63 into a series of basis functions, i.e., core steep ellipsoids (CSEs) which have simple analytical expressions of SL properties such as deflection angles $\bm{\alpha}_{\rm enfw}$ and the lensing potential $\psi_{\rm enfw}$.
@@
they achieved an accuracy of $\sigma=10^{-4}$ in recovering $\kappa_{\rm gNFW}$ using CSEs, with $u_{j}$ spanning a wide range from $10^{-6}$ to $10^{3}$. The amplitude $A_{i}$ and core radius $s_{i}$ are predetermined before evaluating the lensing properties of $\kappa_{\rm gNFW}$ for any given values of $\rho_{s}$, $r_{s}$, and $q$ [^16] Note that $\rho_{s}$ is omitted in Eq. 67 because it acts as a constant scaling factor and does not affect the decomposition process. The corresponding lens potential of an individual CSE is
@@
Appendix C Implementation of the EPL profile
@@
Appendix D Joint modeling with ideal kinematic data across varying source grid resolutions
@@
Appendix E Joint modeling using kinematics data exclude the central bins
(a) No black hole in the composite mass model.
(b) $M_{\rm BH}=7\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ in the composite mass model.
Figure 15: Comparison between joint modeling using full ideal kinematic data (orange contours) and excluding the central regions (green contours). We observe that uncertainties of parameters are enlarged, but the measurements of $D_{\rm d}$ and $\beta_{\rm ani}$ move toward the mock input values using the kinematic map excluding the central regions.
Appendix F The BIC weight factor $f_{\textnormal{BIC}}^{*}$ to joint models
Data Model Source Resolution $\chi^{2}_{\text{dyn}}$ $f^{*}_{\text{BIC}}$ $\Delta\rm BIC$ $\sigma_{\rm BIC}$
@@ Appendix F The BIC weight factor $f_{\textnormal{BIC}}^{*}$ to joint models
COMPOSITE 68 54.11 0.1272 4.154
66 55.08 0.0785 5.120
64 54.61 0.0991 4.654 0.35
$M_{\text{BH}}=1\times 10^{9}M_{\odot}$ 62 54.21 0.1213 4.250
60 54.30 0.1160 4.339
Lensing & Dynamics IDEAL $M_{\text{BH}}=2\times 10^{9}M_{\odot}$ 68 50.61 0.7204 0.6579
66 50.36 0.8171 0.4060
64 50.37 0.8129 0.4163 0.09
62 50.48 0.7712 0.5218
60 50.45 0.7805 0.4977
$M_{\text{BH}}=3\times 10^{9}M_{\odot}$ 68 50.26 0.8612 0.3015
66 49.96 1. 0.0
64 50.00 0.9672 0.0454 0.1
62 50.06 0.9482 0.0989
60 50.03 0.9567 0.0763
$M_{\text{BH}}=4\times 10^{9}M_{\odot}$ 68 50.34 0.8327 0.3676
66 50.10 0.9325 0.1401
64 50.15 0.9065 0.1977 0.08
62 50.20 0.8866 0.2421
60 50.16 0.9050 0.2009
$M_{\text{BH}}=5\times 10^{9}M_{\odot}$ 68 50.87 0.6350 0.9095
66 50.79 0.6581 0.8380
64 50.73 0.6790 0.7756 0.07
62 50.65 0.7080 0.6919
60 50.75 0.6723 0.7953
$M_{\text{BH}}=6\times 10^{9}M_{\odot}$ 68 51.71 0.4162 1.7545
66 51.86 0.3865 1.9026
64 51.81 0.3965 1.8512 0.07
62 51.74 0.4094 1.7875
60 51.67 0.4238 1.7181
Table F.1: continued.
Data Model Source Resolution $\chi^{2}_{\text{dyn}}$ $f^{*}_{\text{BIC}}$ $\Delta\mathrm{BIC}$ $\sigma_{\mathrm{BIC}}$
@@
COMPOSITE 68 52.98 0.2219 3.0197
66 53.55 0.1667 3.5922
64 53.28 0.1904 3.3259 0.19
$M_{\text{BH}}=7\times 10^{9}M_{\odot}$ 62 53.24 0.1949 3.2793
60 53.36 0.1835 3.4002
Lensing & Dynamics IDEAL $M_{\text{BH}}=8\times 10^{9}M_{\odot}$ 68 54.85 0.0870 4.8943
66 55.49 0.0633 5.5294
64 55.11 0.0763 5.1573 0.21
62 55.05 0.0787 5.0941
60 55.13 0.0758 5.1701
$M_{\text{BH}}=9\times 10^{9}M_{\odot}$ 68 56.95 0.0307 6.9940
66 57.94 0.0187 7.9874
64 57.21 0.0269 7.2573 0.34
62 57.22 0.0268 7.2676
60 57.53 0.0230 7.5741
$M_{\text{BH}}=10\times 10^{9}M_{\odot}$ 68 60.93 0.0042 11.06
66 60.67 0.0047 10.71
64 60.73 0.0040 10.98 0.23
62 60.12 0.0052 10.50
60 60.45 0.0052 10.54
EPL Model 68 58.15 0.1338 4.243
66 58.12 0.1356 4.217
64 58.33 0.1227 4.418 0.93
62 60.56 0.0400 6.657
60 58.29 0.1247 4.385